OpenStudy (anonymous):

help!!! Determine two pairs of polar coordinates for the point (4, -4) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

I know it has to do something with the unit circle

OpenStudy (anonymous):

haha, @Mertsj is awesome at explaining this!

OpenStudy (mertsj):



OpenStudy (mertsj):

Find the hypotenuse of that triangle. It will be the radius of your polar coordinate representation.

OpenStudy (anonymous):

how do i do that been awhile

OpenStudy (mertsj):

Pythagorean theorem

OpenStudy (anonymous):

will this give me the answer to the hold problem

OpenStudy (mertsj):

No. Just find the hypotenuse.

OpenStudy (anonymous):

and i got 63

OpenStudy (mertsj):

leave it in exact radical form.

OpenStudy (anonymous):

32

OpenStudy (mertsj):

4^2+4^2=c^2

OpenStudy (anonymous):

16 + 16= 32

OpenStudy (mertsj):

Yes. c^2 is 32 Now find c

OpenStudy (mertsj):

In exact radical form.

OpenStudy (anonymous):

5.6